OpenStudy (jokersmart):

what is the value of the expression when x=9? 5|6-x|

OpenStudy (aihberkhan):

Just plug in 9 for \(x\). So it would become \(5|6-9|\)

OpenStudy (aihberkhan):

Now \(6 - 9\) is \(-3\).

OpenStudy (aihberkhan):

Since it is in the absolute value lines it would become positive \(3\).

OpenStudy (michele_laino):

hint: we can write this: \[\Large \left| {6 - 9} \right| = \left| { - 3} \right| = 3\]

OpenStudy (aihberkhan):

Now do \(5*3\). When you do this you should get \(15\).

OpenStudy (aihberkhan):

That is it! :)

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) If you see that I am online and need help with a question, just tag me in your question! @JokerSmart

OpenStudy (jokersmart):

Oh, thank you @AihberKhan !!! You explain things really well, I'll take note into this. :) I'm currently sick and have to miss school, just finishing my homework. It's bad enough that my teacher doesn't bother to help me- considering I missed her class, lol.

OpenStudy (aihberkhan):

Haha yeah I am not in school either because I am sick as well! Well hope you feel together and glad I helped! :)